Calculate the mole fraction of each component of solution containing 65 g of ethanol (C2H6O) in 350 g of water.



1
Expert's answer
2021-10-27T06:41:23-0400

C = 2/9×65= 14.4/350

H= 6/9×65= 43.33/350

O= 1/9×65= 7.2/350